# YOLO11 and Vision Transformers based 3D Pose Estimation of Immature Green Fruits in Commercial Apple Orchards for Robotic Thinning

### v1

@@ 3.4.2 Depth Anything V2
Field-Level validation of estimated pose
Following the conversion of RGB images to RGB-D format via Vision Transformers, the resultant 3D point clouds were meticulously extracted, as depicted in Figure 7a for the corresponding images. Ground truth values were precisely measured using a vernier caliper, as illustrated in Figure 7b. These in-field measurements served as a benchmark for validating the poses estimated by the YOLO11 deep learning model in 3D image space, where Vision Transformers were employed to transform standard RGB images into detailed RGB-D data without additional depth information from the camera.
The evaluation of the predicted poses against actual field measurements was performed using CloudCompare (Seattle, Washington, USA), a tool for assessing the quality of 3D point clouds. Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) were the primary metrics used to evaluate the accuracy of the pose estimations, specifically measuring discrepancies in the length of green fruits. The RMSE is defined as:
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\text{Predicted Diameter}_{i}-\text{Actual Diameter}_{i})^{2}}$ (5)
@@
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}\left|\text{Predicted Diameter}_{i}-\text{Actual Diameter}_{i}\right|$ (6)
where n is the number of fruits analyzed, Predicted Diameter is the diameter as estimated from the depth maps generated by the DPT and Depth Anything V2 models, and Actual Diameter is the diameter measured in the field using a caliper. These metrics, RMSE and MAE, provided a quantitative measure of the model’s precision in estimating the pose of immature green fruits based on their size, facilitating an objective comparison between the predicted and actual dimensions.
@@ 4 Results and Discussion
4.1 Detection and Pose Estimation Results
The performance metrics of the YOLO11 and YOLOv8 in terms of box detection and pose estimation across all configurations are summarized in Table 1. The results highlight the superior performance of the YOLO11 models, particularly YOLO11n, which achieved the highest precision among all tested configurations. Specifically, YOLO11n recorded a precision score of 0.91, surpassing the best performing YOLOv8 configuration, YOLOv8m, which scored 0.90. Moreover, the inference speed of YOLO11n was noted at 2.7 milliseconds, markedly faster than the quickest YOLOv8 configuration, YOLOv8n, at 7.8 milliseconds.
Table 1: Precision and recall metrics for box and pose estimation for YOLO11 and YOLOv8 configurations. The table focuses on the comparison of these metrics, excluding mAP@50 which is discussed separately.
@@ 4.1 Detection and Pose Estimation Results
In terms of recall, YOLOv8n outperformed all other configurations with a score of 0.905, the highest among its counterparts, while the highest recall within the YOLO11 series was observed for YOLO11x at 0.891. Regarding the mean Average Precision at 50% Intersection over Union (mAP@50), YOLO11s led all configurations with a score of 0.94, closely followed by YOLOv8n, which achieved an mAP@50 of 0.934. These metrics highlight critical aspects of model performance in complex agricultural environments, where not only precision and speed but also recall and mAP@50 are vital for comprehensive performance assessment. The robustness of YOLO11 in detecting and estimating the pose of immature green fruits suggests its applicability across a range of agricultural operations. Potential applications extend beyond robotic thinning to include automated pest management, crop monitoring, yield estimation, and precision irrigation, where the detailed understanding of crop status and dynamics is crucial. These applications demonstrate the broad utility of high-performing detection models in enhancing the efficiency and effectiveness of agricultural practices.
Figures 8a, 8b, and 8c illustrate the comparative effectiveness of the YOLO11 and YOLOv8 models in detection and pose estimation, highlighting their applicability in precision agricultural operations. Specifically, Figure 8a displays the mAP@50 values, whereas Figures 8b and 8c detail the image processing speeds for YOLO11 and YOLOv8 configurations respectively. Given YOLO11n’s superior precision and faster inference speeds, as demonstrated in the processing speed comparisons, it has been chosen as the best-suited model for our robotic thinning project. This decision is scientifically backed by its performance metrics, which promise to enhance efficiency and accuracy in field implementations.
@@
Figure 8: Performance metrics and processing speeds for YOLO11 and YOLOv8 configurations. (a) mAP@50 values comparing both models, (b) Image processing speeds (pre-processing, inference, and post-processing) for YOLO11 configurations, and (c) Image processing speeds (pre-processing, inference, and post-processing) YOLOv8
@@
Furthermore, within the bounding boxes of detected fruits, keypoints were identified by the model, although not all keypoints could be detected when parts of the fruit were not visible, demonstrating a realistic limitation given the partial visibility of some fruits. Such instances are portrayed in Figures 9a and 9b, showcasing the YOLO11n model’s consistent accuracy in detection and pose estimation. Remarkably, the model also identified fruits that were only marginally visible; an example is depicted in Figure 9b, where approximately 10-15% of an immature green fruit at the edge of the image frame was detected. This capability emphasizes the YOLO11 model’s refined sensitivity to minimal visual cues, asserting its robustness in complex agricultural settings.
@@
Moreover, in the depicted results of 10b, the YOLO11n model accurately recognized the object as an immature green apple; however, the pose estimation within the bounding box revealed inaccuracies. Although one keypoint (at the calyx) was correctly identified, the other necessary for a complete pose estimation (at the peduncle) was obscured due to occlusion, a persistent challenge in dense orchard environments. This partial visibility compromises the effectiveness of even the most advanced models like YOLOv10, which demonstrated the highest precision in our study. The upper part of 10b illustrates this limitation: the visible calyx, while detected, does not provide sufficient information to infer the full fruit pose. Future work might benefit from integrating machine learning technologies such as Generative Adversarial Networks (GANs), which have shown potential in reconstructing complete object profiles from limited views Lu et al. (2022). Such approaches could potentially address the occlusions inherent in agricultural settings by extrapolating the unseen portions of fruits, enhancing the accuracy and utility of robotic thinning tools in commercial orchards.
@@
4.2 Evaluation of Visual Transformers in Pose Length Validation
The performance of visual transformers in pose length validation was rigorously assessed by comparing the Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) between estimated and actual measurements of immature green fruit lengths. The validation process involved data from Intel RealSense point clouds, Dense Prediction Transformer (DPT), and Depth Anything models, all cross-referenced against ground truth measurements. Initially, measurements derived from the Intel RealSense exhibited a RMSE of 9.9845 and an MAE of 7.7444, indicating substantial variance from the ground truth. Significant improvements in accuracy were observed with the application of the DPT model, where the RMSE was reduced to 3.2935 and the MAE to 2.6237. The most notable enhancements in pose estimation accuracy were achieved through the Depth Anything model, which demonstrated the lowest errors with a RMSE of 1.5261 and an MAE of 1.2809. These findings underscore the efficacy of using advanced visual transformers like Depth Anything and DPT in enhancing the precision of pose length estimations in agricultural applications, particularly in the context of YOLO11-based fruit detection and pose estimation tasks.
Figure 11: Box plot comparing pose length estimates from different measurement approaches for immature green fruit pose estimation. ’Ground Truth’ represents direct field measurements, ’Intel RealSense’ shows lengths derived from camera-based point clouds, ’DPT’ (Dense Prediction Transformer) and ’Depth Anything’ illustrate lengths from visual transformer-generated depth maps. Variations highlight the disparity in accuracy and reliability of pose estimation across these methods, emphasizing the challenge of occlusion and image quality
@@ 4.3 Discussion
The integration of 3D data has become a cornerstone in the field of robotics Hägele et al. (2016); Amparore et al. (2022), serving as a pivotal element in the advancement of automation technologies. Numerous studies have emphasized that without robust 3D information, effective automation remains unattainable Barazzetti et al. (2010); Fathi et al. (2015); Dadhich et al. (2016); Tian et al. (2020); Xia et al. (2015); Vougioukas (2019); He and Schupp (2018), highlighting the critical role of precise spatial data in enhancing robotic functionalities. The utilization of the Depth Anything V2 transformer, in particular, has proven highly effective in generating detailed 3D mappings from standard RGB images, as illustrated in Figure 13. Subfigure 13a presents the original RGB images captured in a Scifresh orchard, while Subfigure 13b displays the corresponding Intel RealSense point clouds, and Subfigure 13c showcases the Depth Anything V2 generated point clouds. These point clouds were produced at a data collection threshold of 2 feet, demonstrating the transformer’s capacity to accurately interpret and convert 2D images into spatially informative 3D data. This transformation process is crucial for tasks that require precise depth perception, such as robotic fruitlet thinning in densely populated orchards. By providing a clear 3D representation of the environment, Depth Anything V2 facilitates more accurate and efficient robotic interactions, significantly reducing the operational complexity and enhancing the efficacy of automated procedures. The results presented highlight the potential of these advanced vision transformers to not only improve the accuracy of automated operations in agricultural settings but also to drive innovations that could revolutionize the sector.
@@ 5 Conclusion
  • The YOLO11n variant showcased superior precision among all YOLO11 and YOLOv8 models, achieving a precision score of 0.91. In terms of recall, YOLOv8n led with a score of 0.905. Additionally, YOLO11n and YOLOv8x both recorded high mAP@50 values of 0.95, highlighting their effective performance in complex environments.
  • YOLO11 Pose Estimation Performance: Similarly, YOLO11n also excelled in pose estimation precision, matching its detection precision of 0.91, while YOLOv8x displayed the highest recall among the models at 0.91. YOLO11s notably achieved the highest mAP@50 value of 0.94, underscoring its robustness in accurate pose estimation.
  • Image Processing Speeds: YOLO11n was markedly faster in inference speeds, clocking in at 2.7 milliseconds, significantly outpacing the fastest YOLOv8 variant, YOLOv8n, which registered at 7.8 milliseconds. This substantial speed advantage underscores YOLO11n’s suitability for real-time applications in agricultural automation.

### v2

@@ 3.5 Field validation of estimated pose
The evaluation of the predicted fruit size or major axis length against actual field measurements was performed using CloudCompare (Seattle, Washington, USA), a tool for assessing the quality of 3D point clouds. Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) were the primary metrics used to evaluate the accuracy of the pose estimations, specifically measuring discrepancies in the length of green fruits. The RMSE is defined as:
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\text{Predicted Length}_{i}-\text{Actual Length}_{i})^{2}}$ (5)
@@
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}\left|\text{Predicted Length}_{i}-\text{Actual Length}_{i}\right|$ (6)
where n is the number of fruits analyzed, Predicted Length is the major axis length or size of the fruitlet estimated using the depth maps, and Actual Length is the major axis length or size of the fruitlet measured in the field using a caliper. These metrics, RMSE and MAE, provided a quantitative measure of the model’s precision in estimating the pose of immature green fruits based on their size, facilitating an objective comparison between the predicted and actual dimensions.
@@ 4 Results and Discussion
4.1 Fruitlet Detection and Pose Estimation
The comparative assessment of YOLO11 and YOLOv8 models, as presented in Table 1, showed similar performance of YOLO11 and YOLOv8 configurations in terms of precision, with YOLO11n achieving the highest precision of 0.91 and YOLOv8m achieving 0.90. In terms of recall as well, these two configurations performed similarly with YOLOv8n achieving 0.905 and YOLO11x achieving 0.891, demonstrating their strength in capturing all relevant instances. However, the Table 1, showed varying performance of different models in terms of specific performance measures, which emphasizes the importance of selecting the appropriate model configuration based on the specific needs of real-time pose estimation tasks.
Table 1: Precision and recall metrics for box and pose estimations of apple fruitlet achieved with for YOLO11 and YOLOv8 configurations. The table focuses on comparing the models with these metrics excluding mAP@50, which is discussed separately.
@@ 4.1 Fruitlet Detection and Pose Estimation
Figures 7a illustrated mean Average Precision at 50% Intersection over Union (mAP@50) values of all the model configurations tested. As in the case of Precision and Recall, the YOLO11 and YOLOv8 models achieved similar mAP@50), with YOLO11s achieving 0.94 and YOLOv8n slightly behind at 0.934, demonstrating the comparable effectiveness of both models in complex agricultural settings where precision, recall, and mAP@50 are crucial. However, there was a substantial difference in the computational efficiency of YOLO11 and YOLOv8 model configurations, thus appearing as a decisive factor in model selection. Figures 7b and 7c present the image processing speeds, showcasing YOLO11’s faster processing speed compared to the same for YOLOv8 model configurations. Particularly, YOLO11n was the best model configuration in terms of inference speed with 2.7 milliseconds processing time per image, which was significantly better than the quickest YOLOv8 configuration, YOLOv8n, which took 7.8 milliseconds. This considerable difference in inference speeds is pivotal, especially in scenarios demanding rapid data processing such as real-time agricultural operations. The faster processing capabilities of YOLO11n, coupled with its competitive precision and recall, makes it the preferred model for our robotic thinning project. These results not only highlight the model’s robustness in detecting and estimating the pose of immature green fruits but also affirm its potential to significantly enhance operational efficiency and accuracy in the field.
@@
Figure 7: Additional performance metrics for YOLO11 and YOLOv8 configurations; (a) mean Average Precision at 50% mAP@50; (b) Image processing time (pre-processing, inference, and post-processing) for YOLO11 configurations; and (c) Image processing time (pre-processing, inference, and post-processing) for YOLOv8 configurations
@@
Furthermore, within the bounding boxes of detected fruits, keypoints were identified by the model, although not all keypoints could be detected when parts of the fruit were not visible, demonstrating a realistic limitation given the partial visibility of some fruits. Such instances are portrayed in Figures 8a and 8b, showcasing the YOLO11n model’s consistent accuracy in detection and pose estimation. Remarkably, the model also identified fruits that were only marginally visible; an example is depicted in Figure 8b, where approximately 10-15% of an immature green fruit at the edge of the image frame was detected. This capability emphasizes the YOLO11 model’s refined sensitivity to minimal visual cues, asserting its robustness in complex agricultural settings.
@@
Moreover, in the depicted results of 9b, the YOLO11n model accurately recognized the object as an immature green apple; however, the pose estimation within the bounding box revealed inaccuracies. Although one keypoint (at the calyx) was correctly identified, the other necessary for a complete pose estimation (at the peduncle) was obscured due to occlusion, a persistent challenge in dense orchard environments. This partial visibility compromises the effectiveness of even the most advanced models like YOLOv10, which demonstrated the highest precision in our study. The upper part of 9b illustrates this limitation: the visible calyx, while detected, does not provide sufficient information to infer the full fruit pose. Future work might benefit from integrating machine learning technologies such as Generative Adversarial Networks (GANs), which have shown potential in reconstructing complete object profiles from limited views Lu et al. (2022). Such approaches could potentially address the occlusions inherent in agricultural settings by extrapolating the unseen portions of fruits, enhancing the accuracy and utility of robotic thinning tools in commercial orchards.
@@
4.2 Evaluation of Vision Transformers in Pose Length Validation (Fruit’s Major Axis Length)
The assessment of major axis length, an indicator of pose accuracy, was rigorously conducted to evaluate the performance of Vision Transformers in the context of enhanced fruit size estimation. This analysis combined the detection capabilities of YOLO11 with the depth estimation precision of Vision Transformers. The validation process utilized data from Intel RealSense point clouds, the Dense Prediction Transformer (DPT), and the Depth Anything model, all benchmarked against ground truth measurements of immature green fruit lengths. Initially, data from Intel RealSense displayed substantial deviations from actual measurements, with a Root Mean Square Error (RMSE) of 9.9845 and a Mean Absolute Error (MAE) of 7.7444. The application of the DPT model markedly improved accuracy, reducing the RMSE to 3.2935 and the MAE to 2.6237. However, the most significant enhancements in estimating the major axis length were observed with the Depth Anything model, which demonstrated the lowest errors, achieving a RMSE of 1.5261 and an MAE of 1.2809. These results highlight the effectiveness of integrating YOLO11 with advanced visual transformers like DPT and Depth Anything in refining the precision of pose length estimations for agricultural applications, particularly in robotic fruit thinning operations.
Figure 10: Box plot comparing pose length estimates for 70 samples where ’Ground Truth’ represents direct field measurements, ’Intel RealSense’ shows lengths derived from camera-based point clouds, ’DPT’ (Dense Prediction Transformer) and ’Depth Anything’ illustrate lengths from visual transformer-generated depth maps.
Figure 11: Box plot comparing pose length estimates for 70 samples where ’Ground Truth’ represents direct field measurements, ’Intel RealSense’ shows lengths derived from camera-based point clouds, ’DPT’ (Dense Prediction Transformer) and ’Depth Anything’ illustrate lengths from visual transformer-generated depth maps.
@@ 4.3 Discussion
The integration of 3D data is essential in robotics for efficient and effecive navigation and manipulation of target objects Hägele et al. (2016); Amparore et al. (2022). Numerous studies have emphasized that without robust 3D information, effective automation remains unattainable Barazzetti et al. (2010); Fathi et al. (2015); Dadhich et al. (2016); Tian et al. (2020); Xia et al. (2015); Vougioukas (2019); He and Schupp (2018), highlighting the critical role of precise spatial data in enhancing robotic functionalities. The utilization of the Depth Anything V2 transformer, in particular, has proven highly effective in generating detailed 3D mappings from standard RGB images, as illustrated in Figure 13. Subfigure 13a presents the original RGB images captured in a Scifresh orchard, while Subfigure 13b displays the corresponding Intel RealSense point clouds, and Subfigure 13c showcases the Depth Anything V2 generated point clouds. These point clouds were produced at a data collection threshold of 2 feet, demonstrating the transformer’s capacity to accurately interpret and convert 2D images into spatially informative 3D data. This transformation process is crucial for tasks that require precise depth perception, such as robotic fruitlet thinning in densely populated orchards. By providing a clear 3D representation of the environment, Depth Anything V2 facilitates more accurate and efficient robotic interactions, significantly reducing the operational complexity and enhancing the efficacy of automated procedures. The results presented highlight the potential of these advanced vision transformers to not only improve the accuracy of automated operations in agricultural settings but also to drive innovations that could revolutionize the sector.
@@ 5 Conclusion
  • 1.
    The YOLO11n model demonstrated exceptional performance, achieving the highest precision score of 0.91 among all tested configurations and recording one of the top mAP@50 values at 0.95, indicating its effectiveness in complex agricultural settings.
  • 2.
    In pose estimation tasks, YOLO11n matched its high detection precision with a pose estimation precision of 0.91. The YOLO11s variant excelled, achieving the highest mAP@50 at 0.94, which underscores its precision and robustness in accurately estimating poses.
  • 3.
    YOLO11n distinguished itself with superior processing speeds, achieving an inference time of only 2.7 milliseconds, substantially faster than other variants. This rapid processing capability highlights YOLO11n’s suitability for real-time applications in agricultural automation, where efficiency is critical.
  • 4.
@@ 8 Author contributions statement
Our more studies on Sapkota et al. (2024b, a, d, f, c, g, e); Sapkota and Karkee (2024a, b); Churuvija et al. (2025); Sapkota et al. (2025)